OpenStudy (anonymous):

Solve for y. y -1 + - < 15 8

jimthompson5910 (jim_thompson5910):

the inequality is \[\large -1+\frac{y}{8} < 15\] right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

your first step is to move that -1 over, so you add 1 to both sides \[\large -1+\frac{y}{8} < 15\] \[\large -1+\frac{y}{8}+1 < 15+1\] \[\large \frac{y}{8} < 16\] What's your next step?

jimthompson5910 (jim_thompson5910):

you want to isolate y so how do you undo that division by 8?

jimthompson5910 (jim_thompson5910):

do you see what to do?

OpenStudy (anonymous):

I got it, thank you

jimthompson5910 (jim_thompson5910):

ok was just checking